An energy manager is considering three energy saving projects A, B and C that will last up to 5 years. These projects have an initial cost of $2000, $2800 and $3200, respectively. The MARR is 10%. The budget limit on the initial investment is $4800. (i) What type of economic analysis does the energy manager need to perform to make a decision? (ii) List the steps involved in this economic analysis.

(ili) The annual returns from project A, B and C are x+80, x+280 and x+250, respectively, where x is the last 3 digits of your student ID. Carry out the economic analysis for these projects and determine the most preferable combination of the projects The interest table for 1-10% is provided below 1+2+9 12 Marks] Time Value of Money Factors-Discrete Compounding To FindA To Find Given P Given FGiven A Given F Given A Given P Given G Given reo。 03155 23291 | 04762 26210 0.5762 1.2100 13310 0.9366 3.3100 4.0410 6.105101636 0.4021 0229 0.1874 3.1699 4.3781 937430205.4 21.3843 27.975O 0.5132 0.0874 19.4215 33724 7255 0.4241 2.3579 13.5795 15.9374 29.9012 00460 0.0408 0.0357 3.1384 33,4523 14 3.7975 57 3377 13884 4.1772 43.4164 46.5819 49 639S 2.5827 40.5447 .8071 0.121 0.1195 599 51.1591 57 2750 6.9189 71.4027 79 5430 68.4973 0.0140 0.0126 7715 8832 60 6893 63.1402 1403 23 .9543 0.1126 0 1102 25 083 0.0763 71.7773 73.649 121.099 258E-0392372 134.2099 7451E-03 9.3006 148.630 6.726E-03 9.369 164.4940 .079E-03 27 13.1100 28 14 4210 2 15.863 3017.4494 36 30 9127 42 54.7637 497 0172 4171.8719 5.818-03 1.709E 03 58526-04 99418 0304 481 .9137 048 0.1075 0.0630 0.0183537.6370 10E-3.8174 0.0103900.1723 1041E-39.8909 90.5047 1006 940279218 0040 3.284-03 3 035E 03 3 295E-04 857 -04 72 ss.s938 1.046E-035462 03 1.048E-049895 0.1001 99.1419 20 9271E 04 1.079E-05 9 271E 05 1.079E-009.999 00 2.823E 07 3.543-08 2.823-08 3.543-09 10.0000 360 7.9O8E 14 1 2 1 7 966C.15 1 255E-10 10 100.0000 10.000O

student numbers : 283
Answer #1

Answer :

1) We need to calculate present worth of each project and will choose project which have Highest present worth.

2)

Present worth = sum of present worth annual return - initial investment

Since Sum of present worth of annual return =annual return *(1-(1+r)^-n) /r

3)

Annual returns for

A=283+80=$363

B=283+280=$563

C=283+250=$533

Initial investment for

A=$2000

B=$2800

C=$3200

r=10%

n=5 years

Present worth of A=363*(1-(1+10%)^-5)/10% - 2000=-$624.23

Present worth of B=563*(1-(1+10%)^-5)/10% -2800=-$666.23

Present worth of C=533*(1-(1+10%)^-5)/10%-3200=-$1179.93

From above calculation it is clear that project A has highest present worth compare to project B and C.

So Project A will be chosen.since we have budget of $4800 for initial investment and project A has only $2000 as initial investment So we will also choose project B because it's present worth is also greater than Project C and initial investment for project B is only $2800.